Find all points (4,y) that are 10 units from the point (-2, -1).
Answer by stanbon(75887)

Find all points (4,y) that are 10 units from the point (-2, -1).
-------
That would be points (4,y) on a circle with center at (-2,-1) and radius = 10
-------
Circle:: (x+2)^2 + (y+1)^2 = 100
-----
Solve for y if x = 4
-------
6^2 + (y+1)^2 = 100
-------
(y+1)^2 = 64
----
y+1 = 8 or y+1 = -8
----
y = 7 or y = -9
-----------------
Cheers,
Stan H.
--------------